Question 3

Tan Pte Ltd (“TPL”) produces a single product called T01. The product uses material sourced from Europe. The company has sufficient capacity to produce 80,000 units of T01 each month, without the need to increase fixed manufacturing overhead costs. It normally produces and sells 70,000 T01 each month at a selling price of $199 per unit. The company’s unit costs at this level of activity are given below. Fixed manufacturing costs are allocated on the basis of production volume.

Product costs

Per unit ($)

Direct material

100.00

Direct labour

44.50

Variable manufacturing overhead

2.30

Fixed manufacturing overhead

5.00

Variable selling expenses

1.70

Fixed selling expenses

3.50

Total cost per unit

$157.00

Required:

Part A

  1. TPL is considering expansion. The company’s marketing director estimates that sales can be increased by 25% above the current monthly production level. The corresponding increase in fixed expenses is expected to be $803,000 per month. Based on these estimations, should TPL invest in the monthly increase in fixed expenses? Justify your answer by showing the net impact on profit. State any necessary assumption(s) you have to make.

(7 marks)

  1. This month, TPL received a request from an overseas customer for a one-time order of 25,000 units of T01. The import duties incurred by TPL to sell these units overseas will be $2.70 per unit. Another $19,000 is expected to be incurred for permits and licenses to make this overseas sale. The only selling cost for this one-time order will be shipping cost of $2.20 per unit of T01. Calculate the minimum price per unit that TPL should charge for this order. Round your answer to the nearest cent. State any necessary assumption(s) you have to make. Ignore the marketing director’s expansion plan from part (a) for this section.

(8 marks)

  1. Suggest two (2) ways that TPL can increase its factory capacity without outsourcing.

(4 marks)

Part B

This part of the question should be answered independently of Part A.

TPL was just informed that due to a strike in its supplier’s plant, the company is unable to purchase more direct material for the production of T01. The strike is expected to persist for 1 month. TPL currently has enough material on hand to operate at 20% of its normal production level for the coming month.

Alternatively, TPL can close its own manufacturing plant for the month. If TPL’s plant is closed, fixed manufacturing overhead costs would continue at 50%, and fixed selling expenses would be reduced by 30% for the duration of closure.

  1. Would you recommend that TPL close its plant for the month? Show the net impact on TPL’s profit, using incremental analysis, to support your recommendation.

(8 marks)

  1. An outside, competitor manufacturer heard about the strike and has offered to produce and ship directly to TPL’s customers the shortfall in T01 for the coming month. The outside manufacturer will pay for shipping to TPL’s customers, and hence, variable selling expenses incurred by TPL will be reduced to $0.20 per unit for the shortfall units. Based only on the financial information available, compute and explain the maximum unit price TPL should be willing to pay the outside manufacturer. Round your answer to the nearest cent.

(6 marks)

  1. Provide two (2) qualitative factors that TPL should consider when deciding whether to outsource in part (e) above.

(4 marks)
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Tan Pte Ltd (“TPL”)
Variable cost per unit Amount $ Note
Direct Materials              100.00 A
Direct Labor                44.50 B
Variable Manufacturing overhead                   2.30 C
Variable Selling Expenses                   1.70 D
Total Variable cost per unit              148.50 E=A+B+C+D
Sell Price Per unit              199.00 F
Contribution Per unit                50.50 G=F-E
Number of Units         70,000.00 H
Contribution amount 3,535,000.00 I=G*H
Fixed cost
Fixed manufacturing overhead      350,000.00 J= 70,000 units* $ 5 per unit
Fixed selling expenses      245,000.00 K= 70,000 units* $ 3.50 per unit
Total Fixed cost      595,000.00 L=J+K
Net Income 2,940,000.00 M=I-L
Ans 1
Increase in Units         17,500.00 N=H*25%
Contribution Per unit                50.50 G
Contribution Amount      883,750.00 O=N*G
Extra fixed expenses      208,000.00 P= $ 803000- L
Net Income      675,750.00 Q=O-P
As there is net income of $ 675,750 so investment in fixed expenses should be done.
Ans 2- Overseas Market
Total Variable cost per unit              148.50 E
Less: Present Variable Selling Expenses                   1.70 D
Add: Shipping costs                   2.20
Add: Import Duties                   2.70
Revised Variable cost per unit              151.70 R
Additional permits and licenses         19,000.00 S
Number of units         25,000.00 N
Permits and licenses cost per unit                   0.76 T=S/N
Break-even price per unit              152.46 U=R+T
The minimum price per unit that TPL should charge for this order is $ 152.46.
Ans 3 Plant close
Number of units           5,833.33 V=H/12*1
Contribution Per unit                50.50 G
Contribution lost      294,583.33 W=U*G
Savings in Fixed manufacturing overhead by 50%       (14,583.33) X=J/12*1*50%
Savings in Fixed selling expenses by 30%         (6,125.00) Y=K/12*1*30%
Total fixed cost to be avoided      (20,708.33)
Net Financial disadvantage of closing the plant      273,875.00 Z=W+X+Y
Ans 4 20% capacity
Number of units           1,458.33 AA=H/12*1*20%
Contribution Per unit                50.50 G
Contribution earned        73,645.83 AB=AA*G
Fixed manufacturing overhead         29,166.67 AC=J/12*1
Fixed selling expenses         20,416.67 AD=K/12*1
Net Financial advantage at 20% capacity      (24,062.50) AE=AB-AC-AD
So TPL should not close the plant for two months but operate it at 20% capacity.
Ans 5
Variable cost per unit
Direct Materials              100.00 A
Direct Labor                44.50 B
Variable Manufacturing overhead                   2.30 C
Variable Selling Expenses                   0.20 AF
Total avoidable Variable cost per unit              147.00 AG=A+B+C+AF
So the maximum unit price TPL should be willing to pay the outside manufacturer is $ 147 per unit.
Provide two (2) qualitative factors that TPL should consider when deciding whether to outsource in part (e) above.
It should ensure that quality of materials is good and there is no compromise in quality.
It should ensure that the delivery of materials is always in time.